OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

1)find the slope 2) use one points and this formula\[y-y _{E}=m(x-x _{E})\]

OpenStudy (anonymous):

@amirreza1870 is it 7x+y=16?

OpenStudy (anonymous):

\[m=\frac{ 1-2 }{ 5+2 }=\frac{ -1 }{ 7 }\]\[y-2=\frac{ -1 }{ 7 }(x+2)\]7y-14=-x-2 7y+x=-2+14 7y+x=12